How can I write an equivalent expression to 3 2/3 × 8 square root 3 EXPONENT 4

Answer by stanbon(75887)

write an equivalent expression to 3 2/3 × 8 square root 3 EXPONENT 4
---------
(11/3)(8sqrt(3^4)
----
= (11/3)(8*3^2)
----
(11*8*3) = 11*24 = 264
--------------
Cheers,
Stan H.
--------